OpenStudy (anonymous):

Compare the y-intercepts and the rates of change of the following items. The items have the same y-intercept and the same rate of change. The items have different y-intercepts and different rates of change. The rates of change are the same, but the y-intercepts are different. The y-intercepts are the same, but the rates of change are different.

OpenStudy (anonymous):

https://app107.studyisland.com/pics/218789pic28.png

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

The first function is in the form of: y = mx + b Where 'm' is the slope and 'b' is the y-intercept. So what's the y-intercept of y = 0.3x + 3.4?

OpenStudy (anonymous):

How do i find that out?

OpenStudy (igreen):

Compare

OpenStudy (igreen):

y = 0.3x + 3.4 y = mx + b 'm' is slope 'b' is y-intercept Therefore 3.4 is the y-intercept, understand?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

Which of the following relations is not a function? (-4, 4), (0, 2), (9, 1), (-6, 5) (-6, 4), (0, 3), (-4, 1), (9, 2) (9, 4), (-4, 2), (0, 1), (-6, 2) (0, 4), (-4, 2), (0, 1), (-6, 2)

OpenStudy (anonymous):

Which one?A,B,C,or D

OpenStudy (igreen):

Functions can only have 1 x-value.

OpenStudy (igreen):

The x-values in option A are: -4, 0, 9, and -6 Are any of those the same?

OpenStudy (anonymous):

No!

OpenStudy (anonymous):

Which of the following relations is not a function? (-4, 4), (0, 2), (9, 1), (-6, 5) (-6, 4), (0, 3), (-4, 1), (9, 2) (9, 4), (-4, 2), (0, 1), (-6, 2) (0, 4), (-4, 2), (0, 1), (-6, 2)

OpenStudy (anonymous):

Which one ABC or D

OpenStudy (igreen):

Yes, so A is incorrect, because it is a function.

OpenStudy (igreen):

Look at option D, it has x-values of 0, -4, 0, and -6.

OpenStudy (igreen):

Are any of those the same?

OpenStudy (anonymous):

yeah 0

OpenStudy (igreen):

Yes, so that's your answer

OpenStudy (igreen):

Because it has 2 x-values, it is not a function.